# miniCodeProps: a Minimal Benchmark for Proving Code Properties

### v1

@@ Results.
Our baseline results are displayed in Table 3. As expected, our baseline approaches were unable to prove complex function properties such as those in Section 3.4.3. However, failures on many of the comparatively simple Medley properties support the claim that current approaches to neural theorem for mathematics transfer imperfectly to the domain of ITP code verification. However, small measures of success on the termination lemmas indicate that neural theorem proving approaches have the potential to improve on miniCodeProps in the near future.
5 Discussion
Baselines Interpretation
The baselines we present are a reasonable representation of the state of neural theorem proving on code properties at the moment. Current systems can prove the simpler lemmas in the Medley section (especially those with closer ties to math), and are only able to prove the simpler types of lemmas one might use in a larger proof in practice. Proofs of more complex properties requiring understanding larger bodies of code (the Sorting section) remain out of reach for the time being.
Scope
We chose to source miniCodeProps from the TIP dataset to provide some assurance that the code and properties involved were of interest to the broader code verification community. However, we selected a subset of the properties in TIP we judged to be relevant to code, along with termination lemmas, which is a subset of all code properties. Clearly, performing well on this subset is only a proxy for the ability to perform well on all realistic code properties. However, we designed miniCodeProps to target a minimal set of code properties that we would expect a reasonable automated system to be able to prove.
Societal Context
@@ 6 Conclusion
miniCodeProps is intended as a meaningful benchmark for evaluating techniques that automate ITP-based code verification. It contains a diverse array of functions and properties from Tons of Inductive Problems, allowing the possibility of incremental progress in verifying the types of programs contained in TIP. We show that simple baseline approaches from mathematical neural theorem proving currently fall short on much of our benchmark, and hope that miniCodeProps spurs development of future ITP code verification agents.
@@ Appendix A Appendix
have h := numElem_merge_branches_lt Placeholder for specific heaps if needed
have m := merge_elems Placeholder for specific heaps if needed
@@
have h := numElem_merge_branches_lt Placeholder for specific values if needed
have m := merge_elems Placeholder for specific values if needed

### v2

@@ Results.
Our baseline results are displayed in Table 3. All three baselines show nontrivial performance on the easy (medley) properties. This shows that current neural theorem provers have some ability to perform basic inductive proofs in a program verification setting. However, the three methods perform very poorly on the medium and hard properties. The results show that miniCodeProps is sufficient to break current neural theorem proving methods, hence motivating the need for further research. Moreover, it is unlikely that these high-level conclusions would change given more properties drawn from the same distribution, thereby justifying miniCodeProps’ size of 201 properties.
However, it is promising that changing the generation algorithm (namely, to refinement) leads to some additional successes on the medium and hard split. This suggests that further algorithmic development could lead to additional gains. Furthermore, the small ntp-ctx-1.3B model trained on Mathlib shows comparable, or better, performance than GPT-4o. As a result, we view training specialized models for verification-style proofs as a promising direction as well (though requiring nontrivial research due to the scarcity of training data). Based on these observations, we speculate that neural theorem provers have the potential to improve on miniCodeProps in the near future.
Scope.
We chose to source miniCodeProps from the TIP dataset to provide some assurance that the code and properties involved were of interest to the broader code verification community. Clearly, performing well on this subset is only a proxy for the ability to perform well on all realistic code properties. However, we designed miniCodeProps to target a minimal set of code properties that we would expect a reasonable automated system to be able to prove. This is similar to an expert software engineer that spends time working in large, complex codebases, yet is expected to be capable of solving self-contained exercises that target core programming skills.
Societal context.
@@ 6 Conclusion
miniCodeProps is intended as a meaningful benchmark for evaluating techniques that automate ITP-based code verification, including for safe generation of code by AI agents. It contains a diverse array of functions and properties from Tons of Inductive Problems, allowing the possibility of incremental progress in verifying the types of programs contained in TIP. We show that simple baseline approaches from mathematical neural theorem proving currently fall short on much of our benchmark, and hope that miniCodeProps spurs development of future ITP code verification agents.
Acknowledgements
Sean Welleck thanks Convergent Research, the Lean FRO, and the OpenAI Researcher Access Program for their support.
@@ Appendix A Appendix
have h := numElem_merge_branches_lt Placeholder for specific heaps
have m := merge_elems Placeholder for specific heaps
@@
have h := numElem_merge_branches_lt Placeholder for specific values
have m := merge_elems Placeholder for specific values